Continue the two sequence of numbers below and find an equation to each of the sequence
N 1,2,3,4,5,6,7
an 2,5,9,14,20,27,___
bn 1,3,12,60,360,2520,___
Find equation
1
Expert's answer
2021-08-10T11:41:59-0400

The sequences are given as ,

an=2,5,9,14,20,27 .....bn=1,3,12,60,360,2530 ........ 

First sequence is

an=2,5,9,14,20,27 .....


So ,

a1=2 , a2=5 , a3=9 , a4=14

Now ,

a2−a1=3

a3−a2=4

a4−a3=5

·········

an−an-1=n+1


This is cumulative sequence. Therefore by adding we get ,

an−a1=3+4+5+·······+(n+1)

⇒an−2=3+4+5+·······+(n+1)

⇒an=2+3+4+5+·······+(n+1)

⇒an="\frac{n}{2}[2+(n+1)]"

⇒an="\frac{n(n+3)}{2}"


put n=7 ,

a7="\frac{7(7+3)}{2}=\frac{7×10}{2}=35"


Thus , the equation is an="\frac{n(n+3)}{2}"

and a7=35


And second sequence is

bn=1,3,12,60,360,2530 ........

So ,

b1=1 , b2=3 , b3=12 , b4=60

Now ,

"\frac{b_2}{b_1}" =3

"\frac{b_3}{b_2}=4"

"\frac{b_4}{b_3}=5"

······

"\frac{b_n}{b_{(n-1)}}=n+1"

This is multiplicative sequence , Therefore by multiply we get ,


"\frac{b_2}{b_1}×\frac{b_3}{b_2}×\frac{b_4}{b_3}×...........\frac{b_n}{b_{n-1}}=3×4×5.......×(n+1)"


"\frac{b_n}{b_1}=3×4×5×.....×(n+1)"

"\implies\frac{b_n}{1}=3×4×5×.........×(n+1)"

"\implies b_n=3×4×5×.........×(n+1)"

bn="\frac{1×2×3×4×5×......×(n+1)}{2}"


"\implies b_n=\frac{(n+1)!}{2}"


put n=7,

b7="\frac{(7+1)!}{2}=\frac{8!}{2}=\frac{40320}{2}=20160"


Thus , the equation is bn="\frac{(n+1)!}{2}" and b7=20160